40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Feb 15 20:07:28 2013, yuta, Update, Alignment, POP path set up but AS55 is broken 
    Reply  Fri Feb 15 21:22:29 2013, yuta, Update, Electronics, AS55 replaced with POP55 PD 
    Reply  Mon Feb 18 21:43:05 2013, Koji, Update, Alignment, POP path set up but AS55 is broken AS55.pdf
       Reply  Tue Feb 19 00:21:09 2013, yuta, Update, Electronics, POP path set up but AS55 is broken 
          Reply  Tue Feb 19 05:42:28 2013, Koji, Update, Electronics, POP path set up but AS55 is broken 
             Reply  Tue Feb 19 08:06:02 2013, yuta, Update, Electronics, POP path set up but AS55 is broken 
Message ID: 8091     Entry time: Fri Feb 15 20:07:28 2013     Reply to this: 8092   8100
Author: yuta 
Type: Update 
Category: Alignment 
Subject: POP path set up but AS55 is broken 

[Manasa, Yuta]

We set up POP camera and POPDC PD, and centered REFL PDs.
We also tried to center AS55 PD, but AS55 seems to be broken.

What we did:
 1. POP path alignment:
   Shot green laser pointer from ITMX table at where POPDC PD was sitting and centered green beam at optics in the POP path. Steered POPM1/M2 mirrors in the ITMX chamber to make green laser overlap with the PRM-PR2 beam as far as I can reach from ITMX chamber. We removed some ND filters and a BS for attenuating POP beam because POP power was somehow so low. Currently, POP is pick-off of the beam which goes from PRM to PR2.

 2. POP camera and PD:
   We first used camera to find the beam at where POPDC PD was sitting because it is much easier to find focused beam. Put an iris in front of the camera, and put POP DC behind it. Steered a mirror in front of PD to maximize DC output.

 3. REFL PDs:
   Steered mirrors in the REFL path to center the beam and maximized DC outputs, as usual.

 4. AS55:
   AS55 was not responding very much to the flashlight nor AS beam. C1:LSC-ASDC_OUT looked funny. By swapping the ribbon cables of AS55, REFL55, and REFL165, I confirmed that AS55 PD itself is broken. Not the ribbon cable nor PD circuit at LSC rack. I don't know what happened. AS55 was working on Feb 8 (elog #8030).

Result:
  We aligned PRMI coarsely. POP(right above) looks much better than before. REFL (left below) still looks elliptic, but ellipticity differs with the position on the camera. Some astigmatism is happening somewhere. AS (right below) looks pretty nice with MI aligned.


Next:
  1. Fix AS55? Or replace it with POP55 PD, which is currently unused.
  2. Confirm we are getting the right error signals or not, and lock PRMI.

ELOG V3.1.3-